OpenStudy (anonymous):

C2H4+H2<-->C2H6 Kc=9.8*10^-3 Initial concentration of C2H4: Reaction a: 0.1 M Initial concentration of H2: Reaction a: 0.2 M Initial concentration of C2H4: Reaction b:0.5 M Initial concentration of H2: Reaction b: 0.05 M What are the equilibrium concentrations of C2H4, H2, and C2H6?

OpenStudy (aaronq):

why do you have 2 initial concentrations of \(H_2\)?

OpenStudy (anonymous):

It gives reaction a and reaction b for both @aaronq

OpenStudy (aaronq):

oh i see, they're two different scenarios. You need to make an ICE table, and plug your values from the equilibrium line into the equilibrium expression for the reaction and solve it.

OpenStudy (abmon98):

Intial Concentration: 0.1 0.2 - Equilbrium 0.1-x, 0.2-x, x Kc=[C2H6]/[C2H4][H2] 9.8*10^-3=[x]/[0.1-x][0.2-x] Kc is a small number this indicates that more reactants are than products. X is negligible compared to concentration of reactants. [x]=0.1*0.2*9.8*10^-3 x=0.000196 mol/dm^3 you could try out reaction b

OpenStudy (anonymous):

Ok the ones I'm struggling with are the C2H6 equilibrium concentrations @Abmon98

OpenStudy (abmon98):

x represent the concentration of C2H6 at equilbrium